OpenStudy (anonymous):

Resistors of 20ohm,40ohm,50ohm are connected in parallel across a 50V power source.Find the equivalent resistance of the set and the current in each resistor?

OpenStudy (nikvist):

\[R_1=20\Omega\quad,\quad R_2=40\Omega\quad,\quad R_3=50\Omega\]\[\frac{1}{R_e}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}=\frac{1}{10\Omega}\left(\frac{1}{2}+\frac{1}{4}+\frac{1}{5}\right)=\]\[=\frac{1}{10\Omega}\cdot\frac{10+5+4}{20}=\frac{1}{10\Omega}\cdot\frac{19}{20}\quad\Rightarrow\quad R_e=\frac{200}{19}\Omega=10.53\Omega\]\[I_1=50/20 A=2.5A\]\[I_1=50/40 A=1.25A\]\[I_1=50/50 A=1A\]

OpenStudy (shayaan_mustafa):

use resistance combination in parallel. I hope you can do.